OpenStudy (anonymous):

express(sqrt{a}+sqrt{b})div (sqrt{a}-sqrt{b})as a fraction with a rational denomenator

OpenStudy (anonymous):

multiply top and bottom by \[\sqrt{a}+\sqrt{b}\] denominator will be \[a^2-b^2\] and the numerator will be whatever you get

OpenStudy (anonymous):

y nt \[(\sqrt{a}+\sqrt{b})(\sqrt{a}-\sqrt{b})=(\sqrt{a})^2-(\sqrt{b})^2=a-b\]